# Diversity of Transformer Layers: One Aspect of Parameter Scaling Laws

### v1

@@ 1 Introduction
These works suggest that each layer contributes incrementally to the model’s prediction by adding or refining information in the residual stream. However, the relationship between these internal layer-wise mechanisms and the observed scaling behavior of the overall model remains unclear. This disconnect limits efficient performance improvement of Transformers.
@@ 2 Residual Stream in Transformers
where LN is RMSNorm (Zhang and Sennrich, 2019) and is a projection layer. By representing the scaling by RMSNorm as , we can reformulate Equation 1 as follows:
Here, by replacing with , we can reformulate Equation 2 as follows:
We use this equation to investigate the effects of modules in the residual stream theoretically.
@@ 3.1 Prediction Discrepancy
Here, we can reformulate Equation 4 with as follows:
@@ 3.2 Importance of Diversity
First, we focus on the diversity arising from the differences between the predictions of each Transformer layer. From Equations 1 and 6, we can decompose the discrepancy between the Transformer’s predictions and the true logit distribution into bias and diversity terms as in the following theorem.
@@
As shown in Equation 8, MSE can be decomposed into two terms: bias and diversity. Since the diversity term is negative, we can see that when diversity increases, the discrepancy from the true distribution decreases, meaning that prediction accuracy improves. This is different from the bias-variance decomposition, which is widely known in the field of machine learning.
Equations 9 and 10 show the further decomposition of the bias and diversity terms corresponding to the modules on the residual stream in Eq. 1. From this, we can see that each module contributes to both bias and diversity, and that the contribution is weighted by the number of modules. Therefore, we can confirm that MLP layers and attention layers play an important role in Transformers in terms of the number of layers. However, we should not miss the contribution of embedding layers because its importance increases in lightweight models with fewer MLP and attention layers.
@@ 3.3 Limitation of Diversity
By focusing on Equation 11, we can see that the bias and diversity terms share . This suggests that attempting to reduce the bias term may also reduce , which in turn may reduce the diversity term. In other words, Theorem 3 demonstrates the general trade-off relationship between the bias and diversity term, emphasizing the difficulty of maximizing the diversity term without affecting the bias term.
@@ 3.4 Limitation of Stacking Layers
In §3.3, we discussed a trade-off relationship between the bias and diversity terms, and improving the diversity term can potentially improve the prediction performance of Transformers when the bias term is large. A simple method to increase diversity is to add layers. We theoretically investigate the performance improvement and limitations of this method through the following theorems.
@@
where is entropy and is joint mutual information. To minimize the error, we should maximize which can be decomposed to:
where is a set of size and . Note that is interaction information and is conditional interaction information (See Appendix B.4 for the proof.)
@@
In Equation 13, when increases (that means new elements are added to ), Relevancy monotonically increases, whereas Conditional Redundancy and Redundancy do not monotonically increase or decrease. Thus, Information Theoretic Diveristy and do not monotonically increase or decrease, too (See Appendix B.5 for the proof.)
@@
Theorem 6 emphasizes that simply adding layers does not guarantee performance improvement in Transformers. These results seem to indicate that increasing the number of layers in Transformers makes it difficult to improve performance, but at the same time, they suggest that performance can be improved by enhancing diversity. Therefore, as in Theorem 1, diversity is also important from an information-theoretical perspective. However, regarding the information-theoretic diversity, there is a limitation described by the following theorem.
@@
This theorem assumes an ideal situation that the variables in are ideally diverse, which means they are independent given . Here, submodularity refers to the property whereby an increase in input does not lead to an increase in output. In other words, in Transformers, the effect of adding layers on performance decreases as the number of layers increases in the ideal situation. By assuming that the number of layers and the number of parameters are proportional, this result is consistent with the parameter scaling laws, which show that performance converges logarithmically with an increase in the number of parameters. The next theorem reinforces this characteristic.
@@
Supermodularity is a property that the more inputs there are, the greater the risk that additional inputs will result in performance not improving. Therefore, this theorem is consistent with the scaling law, which shows that performance converges logarithmically with an increase in parameters, and reinforces Theorem 7.
Figure 2 shows the example behaviors of non-increasing supermodular and non-decreasing submodular functions. The gradual convergence of values is along with the logarithmic convergence of parameter scaling laws. However, these are based on the conditional independence of the variables in given , which is nearly impossible in the actual situation. In the actual situation, there is a possibility of performance degradation when increasing layers following Theorems 5 and 6. We will check the discrepancy through experiments (§5.5).
@@ 4.1 Difference between probability and logits
Previous discussions have focused on the logits in the output layer of the Transformer. However, when actually using the Transformer, the output is selected from vectors normalized by the softmax layer. Therefore, to investigate whether our theoretical findings are practical, it is desirable to conduct experiments using actual models and datasets. To investigate the behavior of MSE and bias through experiments, it is necessary to prepare the true logit distribution. However, assuming the correct answers in softmax, softmax returns the same value for shifts in the input like , resulting in an infinite number of true logits, which is difficult to handle. To address this issue, we approximate MSE and bias in Equation 8 using the following equation:
@@
4.2 Scaling Law for Model Parameters
Our previous discussion on the parameter scaling laws has been based on the assumption that the number of layers in a Transformer is proportional to the number of parameters. However, this assumption overlooks several important factors. One such factor is the number of dimensions. In many models, the actual number of parameters is represented by the product of the number of dimensions and the number of layers. Therefore, when considering the relationship with the parameter scaling laws in a rigorous manner, it is necessary to take into account elements at the neuron level. Fortunately, it is known that the residual stream in Transformers can be decomposed down to the neuron level (Elhage et al., 2021), allowing us to replace “layers” with “neurons” in our discussion and maintain the validity of our claims.
@@ 5.5 Decrease in Efficiency with Increasing Number of Layers
We verify whether the decrease in performance improvement accompanying the increase in layers shown in Theorems 5, 6, 7, and 8 actually occurs. To do so, we averaged the accuracy and MSE for all tasks in models with no less than 1B of parameters. Figure 6 shows the result. From these results, we can see that the effectiveness of increasing layers actually decreases. Also, the performance degradation caused by increasing layers indicates that the layers are not independent, given answers, which means they are not ideally diverse.
@@ B.4 The proof for Theorem 4
See Zhou and Li [2010] for the proof of the bound and decomposition. Since their claim does not restrict the element of the input and output in their formulation, we can apply their bound and decomposition to the behavior of the residual stream in Transformers.
@@ B.5 The proof for Theorem 5
Conditional Redundancy is a sum of conditional interaction information. We can represent conditional interaction information as follows:
Thus, conditional interaction information can output both positive and negative values. Therefore, it does not satisfy the monotonicity.
Redundancy is a sum of interaction information. We can represent interaction information as follows:
Since interaction information can output both positive and negative values, Redundancy does not fulfill the monotonicity.
@@ B.7 The proof for Theorem 7
Because is a joint mutual information, it is submodular and non-decreasing on under the conditional dependence of the variables in given [Krause and Guestrin, 2005]. According to Zhou and Li [2010], we can convert Information Theoretic Diveristy as follows:
where is conditional total correlation and is total correlation. Here, due to the conditional independence, becomes zero. Since total correlation is a supermodular function [Fujishige, 1978], is a submodular function. Therefore, Information Theoretic Diversity is also a submodular function.
@@ B.8 The proof for Theorem 8
Under the condition, since is a submodular function, is a supermodular function. Only is a term including in both bounds. Therefore, the upper and lower bounds are supermodular functions.

### v2

@@ 1 Introduction
These works suggest that each layer contributes incrementally to the model’s prediction by adding or refining information in the residual stream. However, the relationship between these internal layer-wise mechanisms and the observed scaling behavior of the overall model remains unclear. This disconnect limits the efficient performance improvement of Transformers.
@@ 2 Residual Stream in Transformers
where LN is RMSNorm (Zhang and Sennrich, 2019) and is a projection layer. By representing the scaling by RMSNorm as , we can reformulate Eq. 1 as follows (Chang et al., 2024):
Here, by replacing with , we can reformulate Eq. 2 as follows:
We use this equation to investigate the theoretical effects of modules in the residual stream.
@@ 3.1 Prediction Discrepancy
Here, we can reformulate Eq. 4 with as follows:
@@ 3.2 Importance of Diversity
First, we focus on the diversity arising from the differences between the predictions of each Transformer layer. From Eqs. 1 and 6, we can decompose the discrepancy between the Transformer’s predictions and the true logit distribution into bias and diversity terms as in the following theorem.
@@
As shown in Eq. 8, MSE can be decomposed into two terms: bias and diversity. Since the diversity term is negative, we can see that when diversity increases, the discrepancy from the true distribution decreases, meaning that prediction accuracy improves. This is different from the bias-variance decomposition, which is widely known in the field of machine learning.
Equations 9 and 10 show the further decomposition of the bias and diversity terms corresponding to the modules on the residual stream in Eq. 1. From this, we can see that each module contributes to both bias and diversity, and that the number of modules weighs the contribution. Therefore, we can confirm that MLP layers and attention layers play an important role in Transformers in terms of the number of layers. The contribution of the embedding layer is small based on the number of layers, while in contrast, its relative importance may increase in lightweight models. We empirically investigate these hypotheses through experiments in §5.3.
@@ 3.3 Limitation of Diversity
By focusing on Eq. 11, we can see that the bias and diversity terms share . This suggests that attempting to reduce the bias term may also reduce , which in turn may reduce the diversity term. In other words, Theorem 3 demonstrates the general trade-off relationship between the bias and diversity term, emphasizing the difficulty of maximizing the diversity term without affecting the bias term.
@@ 3.4 Limitation of Stacking Layers
In §3.3, we discussed a trade-off relationship between the bias and diversity terms, and improving the diversity term can potentially improve the prediction performance of Transformers when the bias term is large. A simple method to improve performance is to add layers following the parameter scaling laws and it can also improve the diveristy based on our analysis so far. We theoretically investigate the performance improvement and limitations of this method through the following theorems.
@@
where is entropy and is joint mutual information. To minimize the error, we should maximize which can be decomposed to (Zhou and Li, 2010):
where is total correlation and is conditional total correlation. (See Appendix B.4 for the proof.)
@@
These results seem to indicate that increasing the number of layers in Transformers makes it difficult to improve performance, but at the same time, they suggest that performance can be improved by enhancing diversity corresponding to Information Theoretic Diveristy. Therefore, similar to Theorem 1, diversity is also important from an information-theoretical perspective. However, regarding the information-theoretic diversity, there is a limitation described by the following theorem.
@@
This theorem assumes a kind of lower bound that Conditional Redundancy becomes zero due to the variables in being independent given . Even in this situation, the performance of Transformers increases by adding layers because is non-decreasing. Here, submodularity refers to the property whereby an increase in input leads to decreasing additional benefit in output. Thus, the effect of adding layers on performance decreases as the number of layers increases in this situation. By assuming that the number of layers and the number of parameters are proportional, this result is consistent with the parameter scaling laws, which show that performance converges logarithmically with an increase in the number of parameters. The next theorem reinforces this characteristic.
@@
Supermodularity is a property that the more inputs there are, the greater the risk that additional inputs will result in performance not improving. Therefore, this theorem is also consistent with the parameter scaling laws and reinforces Theorem 7.
Figure 2 shows the example behaviors of non-increasing supermodular and non-decreasing submodular functions. The gradual convergence of values aligns with the logarithmic convergence of parameter scaling laws. However, these are based on a kind of lower bound that Conditional Redundancy becomes zero. In the actual situation, we can expect performance improvement by increased Conditional Redundancy, and it makes the performance improvement non-monotonic, by following Theorems 5 and 6. We check the actual behavior in §5.5.
@@ 4.1 Difference between probability and logits
Previous discussions have focused on the logits in the output layer of the Transformer. However, when actually using the Transformer, the output is selected from vectors normalized by the softmax layer. Therefore, to investigate whether our theoretical findings are practical, it is desirable to conduct experiments using actual models and datasets. To investigate the behavior of MSE and bias through experiments, we should prepare the true logit distribution, . However, assuming the correct answers in softmax, softmax returns the same value for shifts in the input like , resulting in an infinite number of true logits, which is difficult to handle. To address this issue, we approximate MSE and bias in Eq. 8 using the following equation:
@@
4.2 Our Assumption for the Parameter Scaling Laws.
Dimension Size.
Our previous discussion on the parameter scaling laws has been based on the assumption that the number of layers in a Transformer is proportional to the number of parameters. However, this assumption overlooks several important factors. One such factor is the number of dimensions. In many models, the actual number of parameters is represented by the product of the number of dimensions and the number of layers. Therefore, when considering the relationship with the parameter scaling laws in a rigorous manner, it is necessary to take into account elements at the neuron level corresponding to the dimension size. Fortunately, it is known that the residual stream in Transformers can be decomposed down to the neuron level (Elhage et al., 2021), allowing us to replace “layers” with “neurons” in our discussion and maintain the validity of our claims.
Reuse of Layers.
@@ 5.5 Decrease in Efficiency with Increasing Number of Layers
We verify whether the decrease in performance improvement accompanying the increase in layers shown in Theorems 5, 6, 7, and 8 actually occurs. To do so, we averaged the accuracy and MSE for all tasks in models with no less than 1B of parameters. Figure 6 shows the result. From these results, we can see that the effectiveness of increasing layers actually decreases. Also, the performance degradation caused by increasing layers indicates that Conditional Redundancy becomes larger than zero, and the layers depend on each other to predict the answers.
@@ B.4 The proof for Theorem 4
See Zhou and Li [2010] for the proof of the bound and decomposition. Since their claim does not restrict the elements of the input and output in their formulation, we can apply their bound and decomposition to the behavior of the residual stream in Transformers.
@@ B.5 The proof for Theorem 5
Conditional Redundancy is conditional total correlation (conditional multi-information). We can consider its increase when obtains a new element, as follows:
Since is non-negative, Conditional Redundancy monotonically increases.
Redundancy is total correlation (multi-information). We can similarly consider its increase when obtains a new element, as follows:
Since is non-negative, Redundancy monotonically increases. Note that Studeny [2006] introduces the monotonicity of total correlation.
Regarding Information Theoretic Diversity, we can consider its increase when obtains a new element, by utilizing Eqs. 20 and 26 as follows:
When elements in are independent but not independent given , Eqs. 28 is non-negative. On the other hand, when elements in are not independent but independent given , Eq. 28 is non-positive. Thus, Information Theoretic Diversity does not have monotonicity.
Finally, Iyer et al. [2021] indicates the non-monotonicity of mutual information. Therefore, does not satisfy the monotonicity. We can check its increase when obtains a new element, as follows:
When elements in are independent but not independent given , Eq. 29 is non-negative. On the other hand, when elements in are not independent but independent given , Eq. 29 is non-positive.
@@ B.7 The proof for Theorem 7
Because is a joint mutual information, it is submodular and non-decreasing on under the conditional dependence of the variables in given [Krause and Guestrin, 2005]. To check that, we define the change of when adding to as follows:
Since is no less than zero, monotonically increases when increases under the conditional dependence of the variables in given . By utilizing this reformulation,
and utilizing entropy’s submodularity, we can induce the following inequality:
Therefore, is submodular on under the conditional dependence of the variables in given .
Information Theoretic Divergence is a sum of total correlation and conditional total correlation. Both functions are supermodular according to Fujishige [1978], Studeny [2006]. Since is zero under the given condition, Information Theoretic Diveristy becomes
Here is submodular and thus, Information Theoretic Diveristy is a submodular function on .
@@ B.8 The proof for Theorem 8
Under the condition, since is a submodular function, is a supermodular function. Only is a term including in both bounds. Therefore, the upper and lower bounds are supermodular functions and monotonically decrease.